OpenStudy (anonymous):

pls in middle of test help A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate- covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy's bag weighs 1.8 pounds and it cost $1.29. Write an equation in simplest form that represents the problem described here. a. 0.24p + 1.17 = 1.8 b. 0.24p + 1.17 = 1.29 c. 0.24p + 1.602 = 1.29 d. 0.24p + 1.602 = 1.8

OpenStudy (kira_yamato):

0.89 p + 0.65(1.8-p) = 1.29 1.54p = 0.12 OR 0.65 p + 0.89(1.8 - p) = 1.29 1.602 - 0.24 = 1.29